OpenStudy (anonymous):

Find the domain of the function. f(x) = -2x + 4 I'm pretty sure I encountered a problem just like this earlier today...but my mind's still blank.

OpenStudy (jdoe0001):

the domain will be all values "x" can safely take, without throwing off some extraneous value for the function

OpenStudy (jdoe0001):

so one can say in this case, "x" can pretty much take anything you throw at it really, thus \(\large x\in \mathbb{R}\)

OpenStudy (anonymous):

wait...what? o.o

OpenStudy (anonymous):

it can take anything? so are you saying it's \[\infty \]

ganeshie8 (ganeshie8):

\(\large (-\infty , ~\infty )\)

ganeshie8 (ganeshie8):

it can take all the values on number line - from left extreme, all the way to the right extreme

OpenStudy (anonymous):

Thank you both ♥